OpenStudy (mew55):

integral dx/x^3-1? i already got A which is 1/3 but how can i find b and c?

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ x ^{3}-1 } = \int\limits \frac{ A }{ x-1}+\frac{ B }{x ^{2}-1 }+\frac{ C }{ x ^{3}-1 }\] is this your equation?

OpenStudy (mew55):

its \[\int\limits \frac{ dx }{ x^3-1 }\] and the answer from my calculus book shows me with a inverse tangent

OpenStudy (anonymous):

did you setup up A+B+C the same way that I did it?

OpenStudy (anonymous):

FYI, x^3-1= ( x-1)*(x^2+x+1)

OpenStudy (anonymous):

the factorization you have is incorrect, and hence you will not get correct values of B AND C

OpenStudy (anonymous):

you cant factor x^2+x+1

OpenStudy (anonymous):

it does not have real roots

OpenStudy (mew55):

but u have to complete the square right?

OpenStudy (anonymous):

\[\int\limits\ \frac{ dx }{ (x-1)(x ^{2} +x+1)} = \int\limits\ \frac{ A }{ (x-1) }+\frac{ Bx+C }{ (x ^{2}+x+1) }dx\] how does this look?

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

try it out, should work

OpenStudy (anonymous):

ok let's see

OpenStudy (anonymous):

if not try BX^2+Cx+D

OpenStudy (mew55):

i just did that but i am trying to find B and C. i got A as 1/3

OpenStudy (anonymous):

first try with @VeritasVosLiberabit , suggestion, i think you should be able to find b and c, if not try what i suggested with B,C and D

OpenStudy (anonymous):

\[1=(x ^{2}+x+1)A + (x-1)(Bx+C)\] for x = 1 \[1=3A, A=\frac{ 1 }{ 3 }\] \[1=Ax ^{2}+Ax+A+Bx ^{2}+Cx-Bx-C\] \[1=(A+B)x ^{2}+(A-B+C)x+A+C\]

OpenStudy (anonymous):

put x=0 and see what you get

OpenStudy (anonymous):

yeah I think we can solve for C with x=0

OpenStudy (anonymous):

sorry, you need to try , i will verify your work, but not give you exact answer

OpenStudy (anonymous):

\[1=\frac{ 1 }{ 3 }+C, C=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

great now take any value of x, say x=-1 and sub values of A and C to get B

OpenStudy (anonymous):

so with those two values you should be able to get B

OpenStudy (anonymous):

just to let you know A,B and C can be obtained using various methods

OpenStudy (anonymous):

I need to brush up on my factoring

OpenStudy (mew55):

c=-2/3 and b=-1/3. that is wut i got

OpenStudy (anonymous):

does that help

OpenStudy (mew55):

yeah it does XD. thank u

OpenStudy (anonymous):

yw